# Emergent Hierarchy in Localized States of Organic Quantum Chains

### v1

@@ Abstract
keywords:
Organic quantum chains , Carbon nanostructures , Density Functional Theory , Tight-Binding model , Quantum transport , Localized states
[…]
[uff]organization=Instituto de Física, Universidade Federal Fluminense, addressline=Av. Litorânea s/n, city=Niterói, postcode=24210-340, state=RJ, country=Brazil
[…]
[unicamp]organization=Gleb Wataghin Institute of Physics, Department of Applied Physics, State University of Campinas, addressline=Rua Sérgio Buarque de Holanda, 777, city=Campinas, postcode=13083-859, state=SP, country=Brazil
[…]
[unb]organization=Department of Electrical Engineering, College of Technology, University of Brasília, city=Brasília, postcode=70910-900, state=DF, country=Brazil
1 Introduction
@@
Organic Quantum Chains (OQC) are a class of carbon-based systems that exhibit peculiar electronic and structural properties due to their quasi-one-dimensional architecture [1, 2, 3, 4]. These systems are composed of conjugated organic molecules arranged in a chain-like fashion, enabling efficient charge transport and tunable electronic behavior. The properties of OQCs make them promising candidates for applications in molecular electronics, optoelectronics, and quantum-spin computing [5, 6, 7, 4, 8, 9]. Such low-dimensional systems have been synthesized employing self-assembly within chemical routes and direct manipulation with scanning microscope tips [10]. Moreover, when constructed exclusively from carbon-based molecular building blocks (e.g., polycyclic aromatic hydrocarbons, graphene nanoribbons, or conjugated polymers), these systems exhibit emergent electronic localization, topological, and magnetic properties absent in bulk materials [11, 12, 13, 14]. Very recently, nanoscaled chains were synthesized during a growth process of organic molecules on a gold substrate [1].
@@
Figure 1: (a) Experimental STM image highlighting distinct motifs: ring-like structures (red arrow) and chain segments (green arrow) [Adapted [1]]. (b) Molecular precursor used in the synthesis. (c) Schematic representation of the pairing process that connects two precursors. (d) Resulting N1-OQC, representing the smallest unit cell chain segment. (e-m) Optimized geometries of Ni-OQCs for $i=2$ to $10$, illustrating the structural evolution and increasing chain complexity along the periodic direction ($L_{\infty}$).
@@
Stable macrocycle quantum corrals composed of twelve molecular units, marked with a red arrow in Fig. 1(b), are also generated, providing new designed quantum states. Such structures have a precursor unit composed of four hexagons and one pentagon, as depicted in Fig. 1(b). In the sequence, Fig. 1(c) shows the building pair process in the construction of an OQC, exemplified in Fig. 1(d) as an N1-OQC. In this work, we employ various theoretical frameworks for modeling OQCs, with a focus on their electronic structure, quantum transport properties, and potential applications. By combining density functional theory (DFT) with Green’s function (GF) formalism, we provide a detailed understanding of the fundamental properties of OQC and their implications for next-generation molecular devices.
Low-dimensional systems are known to exhibit localized states. A distinct class of localized states, known as compact localized states (CLS), is found in certain low-dimensional periodic systems, such as specific atomic lattices or photonic crystals [11, 15, 16]. Unlike typical wavefunction density probabilities that either fully extend or decay exponentially in space, a CLS has a non-zero amplitude only within a strictly finite and small region, abruptly becoming zero everywhere else. Rather than being caused by randomness or disorder, this extreme form of localization is an intrinsic property of the system’s structure, arising from destructive quantum interference dictated by the specific lattice geometry. The presence of these states is a signature of systems hosting flat bands in their energy spectrum, where the group velocity of particles is zero, effectively trapping them in space and preventing their propagation. Kagomé, Lieb, and fractal lattices, or graphene-based chain lattices [17, 12, 18], are examples of specific geometries that lead to flat bands. The zero-low band dispersion leads to an elevated effective mass of charge holders, obstructing transport and promoting the appearance of topological states of matter, mainly in confined carbon-based structures [19]. Understanding localized states is crucial in designing new devices that harness the robustness and quantum interference of these unique states.
@@
Typically, the coupling of finite molecules to continuum systems results in Fano antiresonance phenomena [20, 21, 22], characterized by a dip in transmission responses due to a destructive interference. On the other hand, in the cases where the molecule states are completely decoupled from the chain system, there is a formation of a perfectly localized mode inside of the continuum state, named a bound quantum state in the continuum (BIC), with no dip on the conductance value related to the decoupled system [23, 24]. BICs are also discussed in wave problems such as acoustics, microwaves and nanophotonics [25, 26, 27]. In this work, we have investigated the emergence of localized states of different natures, creating a unique set of properties that can be directly observed in electronic responses.
2 Theory & Computational details
2.1 Systems Modeling
[…]
@@ 2.1 Systems Modeling
According to the experimental report, the precursor molecule tends to exhibit natural curvature, a feature exploited in the context of electronic confinement in corral-like structures [1]. Based on this behavior, a molecular intercalation strategy was adopted to enable the formation of periodic one-dimensional systems with enhanced geometric stability. In this approach, Ni-OQC is defined as a system whose unit cell contains a pair of $i$ covalently connected molecules arranged in sequence. This configuration promotes mutual compensation of local curvature between adjacent molecular chains of the same length, contributing to the overall structural stability. In the present study, values of $i$ ranging from 1 to 10 were investigated. The number of carbon sites $(N)$ in the unit cell is given by $N=40i$, where $i$ is the OQC index. The corresponding unit cells contained between 60 and 600 atoms (carbon and hydrogen), depending on the number of molecules.
@@
The first stage involved the simultaneous minimization of potential energy and internal forces, excluding thermal effects. The total energy was minimized using the conjugate gradient algorithm. The cell volume was also relaxed under zero pressure in the $z$ direction. To escape from local energy minima, a simulation in the NPT ensemble was performed for 50ps at a temperature of 300K and zero pressure along the longitudinal direction. This step induced realistic structural curvatures along the molecular chains. The system was then linearly cooled from 300K to 1K over 25ps in the same ensemble. An additional simulation of 25ps was carried out at a constant temperature of 1K to eliminate residual thermal fluctuations. Finally, a new energy minimization stage was conducted using the same parameters as in the initial stage to ensure geometric stability before DFT optimization.
Throughout the simulation process, Newton’s equations of motion were integrated using the Velocity-Verlet algorithm [29] with a time step of 0.1fs. Interactions between carbon and hydrogen atoms were described using the Adaptive Intermolecular Reactive Empirical Bond Order (AIREBO) potential [30]. Temperature control was implemented through the Nosé-Hoover thermostat [31, 32].
@@
2.2 Density Functional Theory
@@ 2.3 Tight-binding
Quantum transport properties of OQCs are investigated using the Green function formalism. The Green function $G(E)$ is defined as
$G(E)=\left[EI-H-\Sigma_{L}(E)-\Sigma_{R}(E)\right]^{-1},$ (3)
where $E$ is the energy and $\Sigma_{L}(E)$ and $\Sigma_{R}(E)$ are the self-energies of the left and right electrodes, respectively. The total density of states (DOS) is obtained by $\text{DOS}=-(1/\pi)\text{Im}\{Tr[G(E)]\}$. The transmission function $T(E)$ is given by [44],
$\mathcal{T}(E)=\text{Tr}\left[\Gamma_{L}(E)G(E)\Gamma_{R}(E)G^{\dagger}(E)\right],$ (4)
@@
$T^{\alpha}_{ij}(E)=2\text{Im}\{[G(E)\Gamma_{\alpha}G^{\dagger}(E)]_{ji}H_{ij}\},$ (5)
@@
$t^{\prime}_{ij}=t_{1}e^{-\beta\big(\frac{r_{ij}}{d_{\text{min}}}-1\big)},$ (6)
with $r_{ij}$ being the distance between $i,j$ lattice sites, $t_{1}$ the hopping related to the first nearest-neighbor distance $d_{\text{min}}=\text{1.37}$ Å, and $\beta$ a fitting parameter that controls the range of the hopping energies. As the ratio $r_{ij}/d_{\text{min}}$ is always larger than one beyond the first nearest-neighbors, small values of $\beta$ increase the number of neighbors with non-negligible hopping contributions in the description. The nearest-neighbor distance between atoms inside the unit cell (black bounds in Fig. 1) is within the range 1.37–1.50 Å, obtained from the relaxed structure. This parametrization scheme has been successfully employed for carbon-based mixed geometries in graphene allotropes, such as multidimensional biphenylene systems [46].
@@
where $x$ and $y$ are site coordinates and $\psi_{n}(x,y)$ is the corresponding electronic wave function of the $n^{th}$ state. To compute the delocalized electronic contribution across the system, given by $|\psi_{n}(x,y)|^{2}$, we transform the delta functions into Lorentzian functions, with $\Gamma$ fitted with the experimental results [1].
3 Electronic & Transport Properties
3.1 Organic Quantum Chains (OQCs)
@@
Figure 2: Electronic band structure evolution of the OQC chains for N(1-10). Red curves are the TB results, black and blue the DFT calculations adopting HSE06 and PBE, respectively. The central energy gap is $\Delta E=\text{2.0}$ eV.
Figure 3: (a) Conductance $G(2e^{2}/h)$] and DOS (a.u.) hierarchy of several OQCs for an arbitrary energy range. A pair of plateaus are separated by a gray line. The gray dashed line separates the left and right leads. Plateaus length in eV for (b) Left and (c) Right panels corresponding to the plateaus positioned at left and right of the gray axis, labeled accordingly panel (a). An exponential decaying fit is displayed at the right of panels (b) and (c). Solid black, dashed gray and black, and solid gray lines, respectively, for 1-4 labeled plateaus panel (a).
Building on this structural understanding and based on the energetic analysis detailed in the A, we now examine how the electronic properties evolve along with the $N_{i}$ cases here explored. Fig. 2 shows the evolution of the TB band structures of the system, with an increasing number of precursors inserted in the unit cell of the OQCs. In the first panel ($N_{1}$-OQC), the results are compared between the TB method in red, and DFT, with PBE and HSE06 functionals; blue and black curves, respectively. The best TB parametrization adopted for all $N_{i}$-OQC structures in the present work are the following; $\beta=2.8/d_{\text{min}}=2.04$, $\epsilon=\text{-0.22}$ eV, and $t_{1}=\text{-3.4}$ eV. The exponentially decaying hopping was limited by a cutoff radius of R<4.2R<{4.2} Å.
As expected, the increasing number of precursors in the unit cell (from $N_{1}$ to $N_{10}$) creates new localized states which may be associated with quasi-finite systems. The gap centered at $E=0$ (Fermi energy) is protected against increasing $N_{i}$, remaining the same ($\Delta E\approx\text{2.0}$ eV). This feature is found in both the DFT and TB results. A zigzag-like pattern inside the band structure is observed as the size of the OQC unit cell increases. This effect resembles the folding effect, i.e., when the system is described by a larger unit cell than necessary; however, each of those bands has separate contributions to the electronic properties. The van Hove singularities in the OQCs density of states follow the band structure hierarchy. We further analyze those patterns by choosing a specific energy range where the band separations lead to the formation of mini-bands, as shown in Fig. 2.
The ballistic conductance plateaus shown in Fig. 3(a) are subdivided into compatible sets, compared to the number of precursors inserted into the unit cell. The gray dashed line separates the two groups of plateaus, left and the right, studied with similar subdivision pattern. The hierarchical structure observed in the conductance sequences is also found for other energy ranges, with similar phenomena. The pseudo-gaps around such finite-like states evolve with new states following the hierarchy of increasing number of precursors in the OQC unit cell.
In order to highlight the hierarchical sequence of conductance plateaus, we show in Fig. 3(b) and (c) an analysis between the plateau length (in eV) according to the labels in Fig. 3(a). The both panels (b) and (c) shows a similar trend between the evolution on the left and right features for several OQC sizes following the colored legend for $N_{1}-N_{9}$. In the right side of panels (b) and (c), respectively, we show the exponential behavior of the plateau length (in eV), considering plateaus labeled from 1 to 4, in solid black, dashed gray, dashed black, and solid gray, respectively, for OQCs $N_{1}-N_{9}$. The overall trend shows a slowing exponential decreasing plateau length as the OQC size changes. In particular, the decreasing factor is attenuated when considering the most internal plateaus (labels >2>2). This behavior is evidenced by the exponential decay of the widths associated with these plateaus, which for increasingly larger building blocks indicates a tendency towards electronic localization.
Figure 4: (a) Localization map for the $N_{4}$ OQC chain, plotting charge density ($|\psi|^{2}$) versus site position for all energy states, with detached zoom for energies $E_{a}$, $E_{b}$, $E_{c}$, and $E_{d}$. The lower panel highlights distinct localization behaviors inside the unit cell. (b) Electronic conductance $G$ (in units of $2e^{2}/h$) and density of states (DOS) compared with the electronic bands.
An overall view of the electronic amplitude, $|\psi_{n}(x,y)|^{2}$, gives a fingerprint of the electronic behavior for the energy states and may serve to distinguish between compact localized and spread (or bulk-like) states. The localization pattern is experimentally feasible, as performed for nanographene-based chains [4]. The heatmap in Fig. 4(a) sheds light on the different features of the states by comparing the corresponding renormalized local density (LDOS); purple (yellow) colored range indicates the less (most) probable charge accumulation at the atomic sites displayed in the x-coordinates (Å). For the $N_{4}$ OQC, four energy values were used as examples $E_{a}=\text{-2.18 }$ eV, $E_{b}=\text{-2.20 }$ eV, $E_{c}=\text{6.51 }$ eV, and $E_{d}=\text{6.60 }$ eV] to show the emergence of different kinds of localized states, being conductive like $E_{b}$, or constricted (without transport responses), as $E_{a,c,d}$. This feature was also verified for different OQCs sizes. We see that state $E_{b}$ has a long-spread density probability that crosses the organic chain unit cell. The energy band in this case has a very small dispersion, giving rise to the conductance channel shown in Fig. 4(b). The energies $E_{a}$ and $E_{c,d}$ correspond to localized states at the edges of the unit cell. They do not allow conductance channels. However, $E_{a}$ has localized states spread over a larger range than in the $E_{c,d}$ cases. This can be seen in the lower panel of Fig. 4(a), where the localization tail goes to zero, being localized in fewer sites, with compact behavior, for $E_{c,d}$. Both $E_{a}$ and $E_{c}$ have edge-like localized states, with no percolation between them, and $E_{d}$ has a charge concentration in the center of the unit cell.
The spreading states exhibit an overall uniform charge distribution over the space, as can be seen for several energies in the complete spectrum of Fig. 4(a). These findings show that OQC systems, with their rich electronic behavior, hold significant promise for carbon-based nanodevice applications.
3.2 Coupled Corrals to OQCs
Apart from the isolated organic corral and chains exhibited in the experimental output shown in Fig. 1(a), coupled quantum corrals and chains are also found in different configurations. Here, we investigate the transport and electronic responses of such systems. Most of the structures observed have $N_{1}$ chains coupled to a quantum ring by a single or a small set of precursors. Here we define the system ring + precursor as “molecule”. We start then by comparing the total density of states (DOS) for a single quantum corral and the molecule system, with the reported [1] experimental results of $dI/dV$ curves, as shown in Figs. 5(a,b). The used tight-binding parametrization is the same as discussed before for the chains. We notice that the calculated DOS (gray curves) agree well with the experimental data. Moreover, the quantum corral has an experimental gap $\Delta E=2.35$ eV, while the theoretical model (TB and HSE06 functional) predicts $\Delta E\approx 2.0$ eV. The experimental DOS results exhibit in-gap states (dashed curve in part (a)) due to the gold substrate, absent in the theoretical description.
The energy states of these corral systems are sensitive to atomic coupled structures, as depicted in Fig. 5(b). The energy spectra of the quantum ring, with and without the precursor, are nearly identical, except for the emergence of electronic states within the pseudo-gaps of the corral system, as highlighted by the blue-shaded regions. In contrast, the coupling of different ions at positions P1, P2, and P3 (magenta, green, and yellow disks in Fig. 5(b)) introduces new atomic states not only within the main energy gap (shaded yellow region) but also in other pseudo-gaps, as indicated by the arrows. This sensitivity demonstrates the potential of coupled quantum corrals as nanoscale sensors in advanced device applications.
Figure 5: (a) Left: Experimental dI/dV spectrum of the quantum ring, showing a gap of $\Delta E=2.35$ eV (yellow). Right: Corresponding calculated density of states (DOS). [Adapted from Ref. [1]]. (b) Comparison of electronic properties for the molecule with and without the precursor. The precursor introduces states (blue) inside the original ring’s pseudo-gaps.

### v2

@@ Abstract
I Introduction
@@
Organic Quantum Chains (OQC) are a class of carbon-based systems that exhibit peculiar electronic and structural properties due to their quasi-one-dimensional architecture [1, 2, 3, 4]. These systems are composed of conjugated organic molecules arranged in a chain-like fashion, enabling efficient charge transport and tunable electronic behavior. The properties of OQCs make them promising candidates for applications in molecular electronics, optoelectronics, and quantum-spin computing [5, 6, 7, 4, 8, 9, 10]. Such low-dimensional systems have been synthesized employing self-assembly within chemical routes and direct manipulation with scanning microscope tips [11].
@@
Figure 1: (a) STM image highlighting: ring-like structures (red arrow) and chain segments (green arrow) [Adapted [1]]. (b) Precursor used in the OQCs. (c) Connected pair of precursors. (d) Infinite chain composed of $N_{1}$-OQCs (unit cell marked in gray), and unit cell buildings blocks for $N_{2-9}$-OQCs (rectangular panel) illustrating the structural evolution and chain complexity along the periodic z-direction.(e) Scattering setup for $N_{10}$ with left and right leads.
Moreover, when constructed exclusively from carbon-based molecular building blocks (e.g., polycyclic aromatic hydrocarbons, graphene nanoribbons, or conjugated polymers), these systems exhibit emergent electronic localization, topological, and magnetic properties absent in bulk materials [12, 13, 14, 15]. Very recently, nanoscaled chains were synthesized during a growth process of organic molecules on a gold substrate [1].
@@
Stable macrocycle quantum corrals composed of twelve molecular units, marked with a red arrow in Fig. 1(a), are also generated, providing new designed quantum states. Such structures have a precursor unit composed of four hexagons and one pentagon, as depicted in Fig. 1(b). In the sequence, Fig. 1(c) shows the building pair process in the construction of an OQC, exemplified in Fig. 1(d) as an N1-OQC. In this work, we employ various theoretical frameworks for modeling OQCs, with a focus on their electronic structure, quantum transport properties, and potential applications. By combining DFT with Green’s function formalism, we provide a detailed understanding of the fundamental properties of OQC and their implications for next-generation molecular devices.
Low-dimensional organic systems are characterized by a unique interplay between electron delocalization, driven by $\pi$-orbitals, and spatial confinement arising from the atomic structure of the units. Unlike simple atomic chains, Organic Quantum Chains, such as those reported [1], have large unit cells. The structural complexity gives rise to rich electronic spectra in which dispersive bands that support charge transport coexist with confined molecular states, manifested as highly localized, narrow-band modes. These states arise not from lattice topology, as in Kagomé and Lieb lattices [16, 13, 17, 12, 18, 19], or in higher-dimensional flat band systems [20, 21], but from the specific potential confinement of the molecules raised by an increasing number of precursor units in the unit cell which brings dispersionless states. The electronic probability density in such narrow-band states remains localized within specific regions of the chain.
@@
Typically, the coupling of finite molecules to continuum systems results in Fano antiresonance phenomena [22, 23, 24], characterized by a dip in transmission responses due to destructive interference. On the other hand, in the cases where molecular states are completely decoupled from the chain system, perfectly localized modes are formed inside the continuum state, named a bound quantum state in the continuum (BIC), without a change in the conductance value related to the decoupled system [25, 26]. BICs are also discussed in wave problems such as acoustics, microwaves and nanophotonics [27, 28, 29]. In this work, we have investigated the emergence of localized states of different natures, creating a unique set of properties that can be directly observed in electronic responses. Understanding localized states is crucial in designing new devices that harness the robustness and quantum interference of these unique states.
II Theory & Computational details
II.1 Systems Modeling
To model the OQC formation process observed in the scanning tunneling microscope (STM) experiments [1], we constructed the organic chains by covalently linking $C_{20}H_{12}$ precursors. The formation logic is detailed in Fig. 1 (b-c). Specifically, we removed the two outer hydrogen atoms located at the extremities of the precursor’s longitudinal axis, highlighted with red markers in Fig. 1 (b). This dehydrogenation activates the terminal carbon sites, promoting the C-C coupling that results in the linear chain propagation observed experimentally. The precursor molecule was then replicated along the $z$ axis, maintaining its orientation within the $xz$ plane. The replicated unit underwent sequential rotations about the $x$ and $z$ axes. The molecular units were positioned at an appropriate distance to enable the formation of covalent bonds between the terminal carbon atoms, and a vacuum region of $100\text{\,}\AA$ was added in the $x$ and $y$ directions. The lattice vector along the $z$ direction was adjusted to allow the formation of covalent bonds between the opposite ends of the replicated structure under periodic boundary conditions. The resulting dimer, when infinitely replicated in one direction, constitutes a one-dimensional system here referred to as N1-OQC.
@@ II.1 Systems Modeling
According to the experimental report, the precursor molecule tends to exhibit natural curvature, a feature exploited in the context of electronic confinement in corral-like structures [1]. Based on this behavior, a molecular intercalation strategy was adopted to enable the formation of periodic one-dimensional systems with enhanced geometric stability. In this approach, Ni-OQC is defined as a system whose unit cell contains a pair of $i$ covalently connected molecules arranged in sequence. This configuration promotes mutual compensation of local curvature between adjacent molecular chains of the same length, contributing to the overall structural stability. In the present study, values of $i$ ranging from 1 to 10 were investigated. The number of carbon sites $(\mathcal{N})$ in the unit cell is given by $\mathcal{N}=40i$, where $i$ is the OQC index. The corresponding unit cells contained between 60 and 600 atoms (carbon and hydrogen), depending on the number of molecules.
@@
The first stage involved the simultaneous minimization of potential energy and internal forces, excluding thermal effects. The total energy was minimized using the conjugate gradient algorithm. The cell volume was also relaxed under zero pressure in the $z$ direction. To escape from local energy minima, a simulation in the NPT ensemble was performed for $50\text{\,}\mathrm{p}\mathrm{s}$ at a temperature of $300\text{\,}\mathrm{K}$ and zero pressure along the longitudinal direction. This step induced realistic structural curvatures along the molecular chains. The system was then linearly cooled from $300\text{\,}\mathrm{K}$ to $1\text{\,}\mathrm{K}$ over $25\text{\,}\mathrm{p}\mathrm{s}$ in the same ensemble. An additional simulation of $25\text{\,}\mathrm{p}\mathrm{s}$ was carried out at a constant temperature of $1\text{\,}\mathrm{K}$ to eliminate residual thermal fluctuations. Finally, a new energy minimization stage was conducted using the same parameters as in the initial stage to ensure geometric stability before DFT optimization.
Throughout the simulation process, Newton’s equations of motion were integrated using the Velocity-Verlet algorithm [31] with a time step of $0.1\text{\,}\mathrm{f}\mathrm{s}$. Interactions between carbon and hydrogen atoms were described using the Adaptive Intermolecular Reactive Empirical Bond Order (AIREBO) potential [32]. Temperature control was implemented through the Nosé-Hoover thermostat [33, 34].
@@
II.2 Density Functional Theory
@@ II.3 Tight-binding
Quantum transport properties of OQCs are investigated using the Green function formalism. The Green function $\mathcal{G}(E)$ is defined as
$\mathcal{G}(E)=\left[EI-H-\Sigma_{L}(E)-\Sigma_{R}(E)\right]^{-1},$ (3)
where $E$ is the energy and $\Sigma_{L}(E)$ and $\Sigma_{R}(E)$ are the self-energies of the left and right electrodes, respectively. The total density of states (DOS) is obtained by $\text{DOS}=-(1/\pi)\text{Im}\{Tr[\mathcal{G}(E)]\}$, which is directly related to the summation of Bloch states over all crystal momenta $k$. The transmission function $\mathcal{T}(E)$ is given by [46],
$\mathcal{T}(E)=\text{Tr}\left[\Gamma_{L}(E)\mathcal{G}(E)\Gamma_{R}(E)\mathcal{G}^{\dagger}(E)\right],$ (4)
@@
$T^{\alpha}_{ij}(E)=2\text{Im}\{[\mathcal{G}(E)\Gamma_{\alpha}\mathcal{G}^{\dagger}(E)]_{ji}H_{ij}\}\ .$ (5)
@@
Figure 2: Electronic structure evolution of Organic Quantum Chains. (a) Tight-binding band structures (red curves) for OQC configurations ranging from $N_{1}$ to $N_{10}$. For the $N_{1}$ case, DFT results obtained with the PBE (blue dotted lines) and HSE06 (solid black lines) functionals are also shown, together with the TB parametrization fitted to the HSE06 band structure. (b) DFT-PBE band structures for chains $N_{1}$ to $N_{5}$, confirming the systematic band splitting with increasing unit-cell size while preserving the central energy gap ($\Delta E_{\text{TB}}\approx 2.0$ eV and $\Delta E_{\text{PBE}}\approx 1.30$ eV), in qualitative agreement with the TB results. The TB fitting parameters are $t_{1}=-2.7$ eV, $\varepsilon=-0.55$ eV, and $1.85\leq\beta\leq 2.0$. (c) Top: Zoom-in of the valence-band region for selected chains, highlighting the evolution of the valence-band maximum and the formation of sub-band groups, separated by dashed gray lines. Bottom: Projected LDOS associated with the valence-band maximum, mapped onto the atomic structures, revealing the spatial confinement and hierarchical localization of charge carriers within specific atomic clusters.
Figure 3: Emergence of hierarchical conductance channels in OQCs. (a) Top: Conductance [$G(E)$, solid lines] and DOS (shaded areas) for chains $N_{1}$, $N_{2}$, and $N_{3}$. Black arrows highlight the splitting of a single molecular resonance ($N_{1}$) into NN distinct sub-levels due to inter-precursor hybridization. Bottom: Evolution of the conductance hierarchy for generations $N_{1}$ to $N_{9}$ within the representative mini-band at the valence states (1.6-1.6 eV to 1.0-1.0 eV). Numbered blocks track the splitting of distinct conductance plateaus, divided into “Left” and “Right” spectral regions. (b, c) Evolution of the plateau length (energy width) as a function of the plateau index for the (b) Left and (c) Right spectral regions. The color code corresponds to the $N_{i}$-OQC generation shown in (a). Insets: Exponential decay fits of the plateau length vs. generation index $N_{i}$ for the first four primary plateau families, as shown by the corresponding rectangles, quantifying the spectral densification.
@@
$t_{ij}=t_{1}e^{-\beta\big(\frac{r_{ij}}{d_{\text{min}}}-1\big)},$ (6)
with $r_{ij}$ being the distance between $i,j$ lattice sites, $t_{1}$ the hopping related to the first nearest-neighbor distance $d_{\text{min}}=\text{1.37}$ Å, and $\beta$ a fitting parameter that controls the range of the hopping energies. As the ratio $r_{ij}/d_{\text{min}}$ is always larger than one beyond the first nearest-neighbors, small values of $\beta$ increase the number of neighbors with non-negligible hopping contributions in the description. The nearest-neighbor distance between atoms inside the unit cell (black bounds in Fig. 1) is within the range 1.37–1.50 Å, obtained from the relaxed structure. This parametrization scheme has been successfully employed for carbon-based mixed geometries in graphene allotropes, such as multidimensional biphenylene systems [48, 49].
@@
where $x$ and $y$ are site coordinates and $\psi_{n}(x,y)$ is the corresponding electronic wave function of the $n^{th}$ state. The real-space projection shown in our figures represents the total spectral weight accumulated at site $i$ for energy $E$. To compute the delocalized electronic contribution across the system, given by $|\psi_{n}(x,y)|^{2}$, we transform the delta functions into Lorentzian functions, with $\Gamma$ fitted with the experimental results [1].
III Results
III.1 Organic Quantum Chains
@@
The OQC periodicity along the z-direction ($L_{\infty}$) shown in Fig. 1(d) illustrates the periodic chain for the $N_{1}$ case, where the gray area highlights one choice for the unit cell. The fundamental unit building blocks adopted to construct other OQCs are also illustrated in Fig. 1(d) at the detached rectangle. The optimized structures obtained allow us to model not only the connectivity but also the structural evolution of the OQCs’ curvature. For $N_{2}$-N9, the translational unit cell contains the entire sinusoidal segment, capturing the strain compensation mechanisms that stabilize the complex organic nanostructures. The scattering setup adopted for the transport analysis of the OQCs is depicted in Fig. 1(e) for the $N_{10}$-OQC case. The central region of electronic scattering is connected with the semi-infinite left and right leads.
Building on the structural understanding and based on the energetic analysis detailed in the Supplementary Material (SM) [50], we now examine how the electronic properties evolve along with the $N_{i}$ cases here explored.
The calculated band structures shown in Fig. 2 reveal the intricate evolution of the electronic spectrum as the complexity of the unit cell increases from $N_{1}$ to $N_{10}$, with length $a$ according to the OQCs. In the panel corresponding to N1-OQC, we present first-principles results obtained using both the DFT-PBE functional (blue dots) and the hybrid HSE06 functional (black lines). The TB parameters were chosen to accurately reproduce the upper valence and lower conduction bands obtained with the HSE06 potential. The entire sequence of TB band structures for larger unit cells (red curves) was then generated using the same parameter set. The best TB parametrization adopted for all $N_{i}$-OQC structures in the present work are the following; $\beta=2.8/d_{\text{min}}=2.04$, $\epsilon=-0.22$ eV, and $t_{1}=-3.4$ eV. The exponentially decaying hopping was limited by a cutoff radius of R<4.2R<{4.2} Å. To validate the reliability of the TB model for extended systems with increasingly complex unit cells, we performed additional DFT-PBE calculations for configurations ranging from $N_{1}$- to $N_{5}$-OQC, as shown in Fig. 2(b). These results demonstrate that the TB parameterization consistently captures the essential physics as chain length increases, providing an efficient and accurate alternative when the computational cost of DFT calculations with complex functionals, such as HSE06, renders them practical. Both approaches reproduce the characteristic splitting of molecular levels and the persistence of the central energy gap ($\Delta E\approx 2.0$ eV, as obtained from the HSE06 fit).
As the number of precursors inside the unit cell grows, the electronic bands change accordingly by systematically subdividing into a set of mini-band spectra following the discrete molecular orbitals of the individual building blocks. A detailed inspection of the valence bands, shown in the upper panel of Fig. 2 (c), reveals the transition of dispersive states to flat bands as the OQC changes from $N_{3}$- to $N_{9}$-OQC, when considering, for instance, the valence band maximum (VBM), highlighted in green. To illustrate the physical features of these flat bands, we compute the real-space-projected LDOS of the VBM of the proposed chains. As displayed at the lower part of Fig. 2 (c), the electronic probability densities are not always uniformly distributed along the chain. Depending on the $N_{i}$-OQC length, they may exhibit spatial confinement at specific segments of the precursor units. In the case of dispersive bands, the electronic state is delocalized over the entire unit cell, like $N_{3}$-, $N_{4}$-, and $N_{5}$-OQC cases, while flat bands lead to states localized at atomic clusters, as observed for $N_{6}$-, $N_{7}$-, and $N_{9}$-OQC, with no percolation within the entire unit cell.
To further understand the electronic evolution of the OQCs, we have analyzed the transport signatures of the valence states. Fig. 3(a) illustrates the formation of a hierarchy of conductance plateaus within a representative mini-band as previously discussed in Fig. 2(c) with energy range 1.6-1.6 to 1.0-1.0 eV. We notice that the outermost plateaus, labeled as 1 at left and right, respectively, have a decreasing length as the system grows, with a hierarchy induced by the subdivision of the bands being proportional to the number of inserted precursors pairs in the unit cell. The physical origin of this hierarchy is the hybridization of the discrete states molecular orbitals from the precursor units. As shown in the top panels of Fig. 3(a), a single transmission plateau observed in $N_{1}$ splits into exactly two sub-channels in the $N_{2}$, and three in the $N_{3}$ case. This pattern generalizes for any $N_{i}$ chain. The lower panel of Fig. 3(a) tracks this evolution up to $N_{9}$.
The emergent hierarchy does not arise from a stochastic increase in energy levels, but from a well-defined, structured subdivision of the spectrum. The process follows a self-similar pattern across different energy ranges, as exhibited in Fig. 3 (a) with a systematic partitioning quantified in panels (b) and (c) of Fig. 3. By analyzing the evolution of plateau length for the first 3 families, defined as the energy width of a single transmission window, we observe a robust exponential decay with rate parameter $b$. The mathematical trend provides a quantitative parameter for the emergent hierarchy in such systems. Moreover, it reflects a regular pattern of spectral subdivision where the fixed energy bandwidth of the initial $N_{1}$-OQC resonance is iteratively subdivided into $N$ discrete sub-levels and can be predicted as the chain length grows. These findings show that OQC systems, with their rich electronic behavior, hold significant promise for carbon-based nanodevice applications.
III.2 Coupled Corrals to OQCs
Apart from the isolated organic corral and chains exhibited in the experimental output shown in Fig. 1(a), coupled quantum corrals and chains are also found in different configurations. Here, we investigate the transport and electronic responses of such systems. Most of the observed nanostructures have $N_{1}$ chains coupled to a quantum ring by a single or a small set of precursors. We start by analyzing the total DOS of a single quantum corral as reported in the $dI/dV$ experimental results [1], as shown in Figs. 4(a). The used tight-binding parametrization is the same as discussed before for the chains. The experimental results show a fundamental gap of $\Delta E\approx 2.35$ eV (yellow region), while the TB model predicts a slightly smaller gap of $\Delta E\approx 2.0$ eV. This underestimation is typical for DFT-based parameterizations [51]. Notably, the experimental spectrum exhibits a broad mid-gap feature (dashed gray line) attributed to substrate hybridization. As our isolated model does not include the gold substrate, we investigate similar in-gap phenomena arising from local chemical hybridization by considering coupled ions at specific sites (P1, P2, P3) as shown in Fig. 4(b). The broken molecular symmetry introduces localized states deep within the fundamental gap (near $E=0$) and inside the valence pseudo-gaps. The discrete impurity levels (colored markers) effectively provide accessible quantum states which are often observed in substrate-coupled or doped molecular corrals. Features observed within the gap are attributed to Au substrate states rather than intrinsic molecular orbitals, being consistent with previous reports [1, 52, 10]. We conclude that the sensitivity of the rings to absorbed ions demonstrates the potential application within nanoscale sensors.
Figure 4: (a) Left: Experimental dI/dV spectrum of the quantum ring, exhibiting a fundamental gap of $\Delta E\approx 2.35$ eV (yellow shaded region) [Adapted [1]]. Right: Corresponding theoretical DOS for the pristine ring. (b) Left: Structural model showing the positions of coupled ions P1 (orange), P2 (magenta), and P3 (blue). Right: Discrete energy spectrum comparing the pristine ring states (gray circles) with the modified states induced by the ions. The colored markers explicitly show the emergence of localized in-gap states near the Fermi level and within the valence pseudo-gaps.